Required information Problem 10-74 (LO 10-4) [The following information applies to the questions displayed below.] Bethany incurred $38,000 in research and experimental costs for developing a specialized product during July of year 1. Bethany went through a lot of trouble and spent $22,000 in legal fees to receive a patent for the product in August of year 3. Bethany expects the patent to have a remaining useful life of 10 years. (Do not round intermediate calculations. Round your final answers to nearest whole dollar amount.) Problem 10-74 Part-a a. What amount of research and experimental expenses for year 1, year 2, and year 3 may Bethany deduct if she elects to amortize the expenses over 60 months? Research and experimental expenses Year 2 3
Answer #1
Year Research and experimental expenses
1 3800 =38000/60*6
2 7600 =38000/60*12
3 4433 =38000/60*7
Note:
In year 1, there will be 6 months of amortization (July-Dec)
In year 2, there will be 12 months of amortization (Jan-Dec)
In year 3, there will be 7 months of amortization (Jan-July)